# On the polarimetric response of the Nançay Radio Telescope and its impact on precision pulsar timing

### v1

@@ 2 Does the polarimetric response of the NRT depend on the observed direction?
Initial results from the analysis of the rotating horn observations made between MJDs 59368 and 59373 revealed that the “0/0” model, i.e., the simplest model in which the differential gain and phase do not depend on declination and hour angle, gave the best results in terms of the test statistic we chose for comparing results for the different configurations: a “quasi-AIC” (hereafter qAIC), consisting in an average value of Akaike Information Criterion (AIC; Akaike, 1973) scores over the frequency channels in which the fit was successful. A number of metrics can be used for comparing results for the various configurations, in these cases where the number of degrees of freedom varies from one configuration to another: for example, the AIC, the Bayesian Information Criterion (BIC; Schwarz, 1978), the Geometric Information Criterion (GIC; Kanatani, 1998), or the Stochastic Information Complexity (SIC; Rissanen, 1989). However, in our pulsar observations declinations are not uniformly sampled and only different declination values are tested, where is the number of pulsars observed, whereas hour angles are sampled uniformly and the number of hour angles tested is much larger than . As a consequence, strong asymmetries exist between the two dimensions, in terms of the information available to constrain model parameters. The AIC, BIC and GIC metrics, that are in principle designed to penalize model complexity by accounting for the varying numbers of degrees of freedom in the compared models, cannot account for such strong asymmetries in the sampled dimensions. The SIC metric, which we also tested, uses the logarithm of the determinant of the Hessian metrix to penalize model complexity. However, we found that the Hessian tends to be ill-conditioned in the cases where the polynomial order in declination approaches the number of pulsars available to constrain any model of variation along this dimension. Finally, additional complexity emerged from the fact that pcm analyses tended to successfully fit the available data in fewer frequency channels when polynomial orders increased. With all the above-mentioned caveats in mind, we chose to adopt the simple AIC metric in our tests, but dividing AIC scores by the number of frequency channels with successful fits, to form the “quasi-AIC” metric. In future work we will further investigate the issue of the choice of metric for comparing the results from these analyses. We note, also, that in addition to qAIC scores, we used pulsar observation S/N values as another metric for comparing the different configurations, as will be presented later in this section.
@@
In Fig. 4, we show the best-fit calibration parameters and Stokes parameters of the noise diode from the “0/0” analysis of the rotating horn observations made between MJDs 59368 and 59373, and those from the analysis of the observation of J07422822 on MJD 56368. In general the results from the two analyses are very similar. In some cases, such as those of the differential gain or the ellipticity parameters and (assumed to be equal), some small differences are found. Nevertheless, best-fit parameters from the “0/0” analysis tend to present smoother variations across the bandwidth than those from the analysis of J07422822 and preceding noise diode observation, with less pronounced variability and fewer outliers (see the best-fit values for the orientation parameter for a particularly clear example). This suggests that the instrumental response inferred from the “0/0” analysis better represents the actual response than that inferred from the analysis of J07422822, and highlights the usefulness of jointly analyzing multiple pulsar observations at once.
@@
Most other analyses resulted in higher qAIC scores than the simplest “0/0” model, with the exceptions of the “0/1”, “1/0” and “1/1” configurations which gave very similar qAIC scores. In order to discriminate between these models, we conducted two other series of rotating horn observations of pulsars at different declinations, between MJDs 60240 (23 October 2023) and 60242 (25 October 2023), and on MJD 60461 (31 May 2024). These second and third observation sessions are described in Table 1. The pulsars observed during these two new sessions were the same as in the first, with the exception of PSR J0358+5413 (B0355+54, Manchester et al., 1972) which was added to the target list for the second session. Beyond the fact of simply taking more data, we took the opportunity of these new observing sessions to observe over longer durations than what is typically done with the NRT, and thereby observe over wider HA ranges, close to the limit that can explored with this telescope. With these longer observations, we wanted to increase our sensitivity to potential dependencies of calibration parameters on HA.
@@
We carried out pcm analyses for each of these three observation sessions, testing the same combinations of 2D polynomial orders as described above. The results from these combined analyses are summarized in Table 2. In this table, the “Total AIC” column gives the sum of the AIC scores from the three observation sessions, the “Total nchan” columns gives the total number of frequency channels in which the fit was successful (since the NUPPI backend splits the 512 MHz of bandwidth into 128 channels of 4 MHz each, the maximum possible value is ), and the “qAIC” column gives the quasi-AIC scores, i.e., the total AIC scores divided by the total number of frequency channels with a succesful fit. As can be seen from the table, for configurations with polynomial orders in declination larger than three, the total numbers of frequency channels with a successful fit are smaller than under the simplest configurations, and even much smaller in some cases (see e.g. the results for the “6/3” analysis), indicating unstable fits. Quasi-AIC scores are also higher, with the exception of the “4/3” configuration which gave the lowest qAIC score among all analyses. In this analysis, however, successful fits were found in much fewer frequency channels than for the simplest analyses, so that we did not consider the “4/3” configuration further. In fact, the simplest configurations with polynomial orders in declination of up to two are found to produce the best results, both in terms of qAIC score and number of frequency channels with successful fits.
, or each analysis configuration. The following four columns give the same information, sorted by qAIC score.
@@
Taken at face value, Table 2 suggests that “0/1”, “0/2”, “0/3” and “1/1” configurations produce better results than the simplest model, the “0/0” configuration. To test whether these seemingly better configurations also constitute better calibration solutions for normal mode observations (i.e., regular NRT observations of pulsars with no horn rotation), we analyzed selections of NUPPI observations of MSPs, calibrating the observations with the parameters found by pcm for each of the three sessions and configurations. The selections consisted in MSP observations conducted within 30 days of MJDs 59370, 60240, and 60461, i.e., close in time to the three special mode observation sessions presented in Table 1. We selected normal mode MSP observations with S/N values larger than 100, to limit the sample of observations to analyze222The S/N values considered here were those obtained after calibrating the NUPPI data with the method presented in Guillemot et al. (2023), i.e., using calibration solutions derived from rotating horn observations of PSR J07422822.. In this work, S/N values were calculated using Eq. (7.1) of Lorimer & Kramer (2004), that is:
@@
where denotes the amplitude of the profile bin, and represent the off-pulse standard deviation and mean amplitude, and is the equivalent width, expressed in number of profile bins, of a top-hat pulse with the same area and peak height as the profile. After the above-mentioned selection cuts, we obtained datasets comprising 52, 55 and 57 MSP observations, respectively.
@@
In Fig. 5, we compare the S/N values of the total intensity pulse profiles for the three MSP datasets under different calibration methods. In the left-hand panel we compare S/N values as obtained when calibrating MSP observations with the “0/0” model, and those obtained using the IFA calibration model. The improved calibration produces increased S/N values in more than 75% of the considered observations, in line with the results of Guillemot et al. (2023). The middle and right-hand panels show comparisons of the “0/0” model S/N values with those obtained with the “0/1” and “1/1” calibration models. S/N values obtained with the “0/0” and “0/1” models are very similar, with differences very close to 0 in most cases. Yet, more than 60% of the S/N values are larger with the “0/0” model than with the “0/1” model. As can be seen from the right-hand panel, S/N values obtained when calibrating the data with the “1/1” model (i.e., with differential gain and phase values that depend on declination) are degraded in almost 80% of cases, with, in particular, large differences at high declination values. We also calibrated the same datasets using “0/2” and “0/3” models, finding that these models degraded S/N values in 79% and 78% of the cases, respectively. We therefore find that the “0/0” model, in which the differential gain and phase parameters do not vary with declination and hour angle, produce better results than the other models on normal mode MSP observations, in terms of S/N values. This suggests that the better qAIC values obtained for the “0/1”, “0/2”, “0/3” and “1/1” configurations (see Table 2) resulted from overfitting of the data, and not from actual improvements of the models. Finally, we repeated the same analyses as those presented earlier in this section, this time allowing the feed orientation parameter and the feed ellipticities to vary with hour angle and declination using the same parameterization as presented in Eq. 3, and fitting constant values for the differential gain and phase parameters. The conclusions from this analysis were similar: better qAIC scores were found for a few of the tested configurations (the same configurations as in Table 2) than for the “0/0” one. However, the application of these calibration solutions to normal-mode observations of MSPs made within 30 days of MJDs 59370, 60240, and 60461 again led to degraded S/N values. We therefore take our results to mean that the polarimetric response of the NRT does not seem to depend on hour angle and declination.
@@ 3 Improving polarization calibration over the entire NUPPI dataset
The first analysis step was thus to construct a high S/N well-calibrated polarimetric profile, of a particular pulsar. Ideally, the chosen pulsar should be bright enough for METM to be able to determine accurate calibration solutions for the individual 4-MHz frequency channels. Tests we conducted showed that this condition requires individual observations to have S/N values of several thousands at least. Also, there needs to be numerous observations of the desired pulsar, between August 2011 and the end of the considered NUPPI dataset. Two pulsars stand out, in the case of the NUPPI data archive: J0953+0755 (B0950+08), and J1136+1551 (B1133+16; Pilkington et al., 1968). Among these two objects, J0953+0755 has the largest flux density at 1.4 GHz: against mJy (Jankowski et al., 2018), and is therefore typically detected with higher S/N values with the NRT. J1136+1551, however, has been observed much more often than J0953+0755 with NUPPI: almost four times more often, so that the density of J1136+1551 observations is larger. Given that both pulsars appeared to be suitable choices for the METM analysis, we conducted analysis tests using both of them.
@@
We selected individual 1.4 GHz NUPPI observations of J0953+0755 and J1136+1551, conducted close in time to the first session described in Table 1 and with high S/N values. The selection led us to choose the MJD 59372 observation of J0953+0755 listed in Table 1, and a normal mode observation of J1136+1551 conducted on MJD 59376 (11 June 2021). The two observations were calibrated using a modified version of the best-fit calibration solution obtained from the “0/0” analysis of the rotating horn observations made between 59368 and 59373. This modified version consisted in a version of the best-fit calibration solution in which parameters for the frequency channels with no data (e.g., due to RFI mitigation) were linearly interpolated using data from the neighboring channels. The selected and calibrated NUPPI observations were integrated in time, and the full frequency resolution of 128 channels was kept. Both observations had high S/N values: 14000 for J0953+0755 and 4480 for J1136+1551.
@@
All other 1.4 GHz NUPPI observations of J0953+0755 and J1136+1551 (and preceding observations of the noise diode) were then compared to the above-described reference polarimetric profiles, using the METM method of pcm. In addition to determining the gain , differential gain , differential phase and the Stokes parameters of the reference noise diode, pcm fits for four additional parameters: the difference in the receiver ellipticities, , the sum of the ellipticities, , the difference in the receiver orientations, and the sum of the orientations, . Inspection of the analysis results for the individual observations of J0953+0755 and J1136+1551 enabled us to observe that the best-fit calibration parameters as determined from analyses of J1136+1551 observations displayed very strong variability across the bandwidth, strong variability from epoch to epoch, and large numbers of outlier values. On the other hand, best-fit results from the analysis of J0953+0755 observations displayed much smoother variations, and with low numbers of outlier values. The much better fits obtained in the case of J0953+0755 led us to choose this pulsar for the subsequent analyses, and discard J1136+1551.
@@
In order to calibrate a normal mode NUPPI observation of a pulsar at a given epoch, one possibility is to use the calibration parameters derived using the METM method from the J0953+0755 observation made closest to that epoch. However, in further investigating the results for individual J0953+0755 observations, we noticed that the frequency variations of the best-fit parameters were consistent over time segments. More specifically, in each of these time segments, the individual parameters appeared to be scaled and/or offset versions of common functions, to be determined. Instead of calibrating pulsar observations using the calibration parameters derived from the J0953+0755 observation conducted closest in time, we thus chose to explore a different strategy: determine the boundaries of the above-mentioned time segments, determine the “archetypal” functions describing the frequency variations of the different calibration parameters over those time segments, and eventually use this information to find the calibration parameters at the considered epoch.
@@
As mentioned above, best-fit results from the analysis of J0953+0755 were more suitable for the analysis presented here than those obtained from the analysis of J1136+1551 observations. Nevertheless, the larger density of J1136+1551 observations made the data for the latter pulsar more appropriate for finding the boundaries of the different time segments. In Fig. 6 we show three of the calibration parameters determined with the METM method as a function of time, for PSR J1136+1551, and for the 64th 4-MHz frequency channel. As can be noted from this figure, NUPPI observations of J1136+1551 have been conducted regularly over the entire dataset, with a small number of gaps that generally coincide with time intervals over which the NRT was inactive (See Sect. 3.1 of Guillemot et al., 2023, for a list of the main interruptions between 2011 and 2023). Vertical black lines indicate J0953+0755 observations, also conducted regularly since August 2011, but with smaller cadence. We visually inspected METM results for J1136+1551, searching for time discontinuities in the frequency variations of best-fit parameters. In Fig. 7 we illustrate, with the example of the absolute gain parameter , the variations seen from one time interval to another in the evolution of calibration parameters with frequency. This step enabled us to determine the boundaries of the above-mentioned time segments, indicated as dashed vertical lines in the figure.
@@
In the following step, for each time segment we selected a single observation and used the calibration parameters for this observation as initial guesses of the “archetype” functions for the corresponding time segment. The single observations we chose as starting points had good fits and low numbers of excised frequency channels. The frequency evolution of the parameters for the chosen observations were spline smoothed to form the initial guesses of the archetypes. Then, the parameters for the other observations in the time segments were fitted to their corresponding archetypes. The gain parameter was fitted using the following function:
@@
where represents the value of the archetype function at frequency , is the value of the parameter to be fitted and is a scale parameter. The and parameters, found to be close to zero in most channels and observations, were fitted to the following function:
@@
where is an offset parameter. Finally, the other parameters (, , , , and Stokes parameters , , and for the reference noise diode) were fitted as follows:
@@
In the following analysis step we refined the archetypes for each parameter and time segment, by forming weighted averages of the calibration parameters within each segment, correcting for the scale and offset parameters and for each observation determined from the previous step. This analysis step enabled us to obtain improved parameter archetypes, that in some cases included values for more frequency channels, i.e. the frequency channels that were absent in the chosen starting points but present in other observations from the same time segment. The improved parameter archetypes were then used to determine new estimates of the scale and offset parameters and using Eq. 7 to 9, for each calibration parameter and time segment. In Fig. 8 we plot the values of the calibration parameters as a function of frequency for two of the time segments. Parameter values were multiplied and/or offset using the and factors determined from this iteration of the analysis. The scaled and offset calibration parameters match the corresponding archetypes well, validating the construction of the archetypes and the measurement of the and parameters for each observation. Fig. 8 illustrates the fact that calibration parameters have common frequency evolution properties in each time segment, and that the frequency dependence of the parameters change significantly in other time segments.

### v2

@@ 2 Polarimetric response of the NRT: Possible dependence on the observed direction
Initial results from the analysis of the rotating horn observations made between MJDs 59368 and 59373 revealed that the 0/0 model, i.e., the simplest model in which the differential gain and phase do not depend on declination and hour angle, gave the best results in terms of the test statistic we chose for comparing results for the different configurations: a “quasi-AIC” (hereafter qAIC), consisting in an average value of Akaike Information Criterion (AIC; Akaike, 1973) scores over the frequency channels in which the fit was successful. A number of metrics can be used for comparing results for the various configurations, in these cases where the number of degrees of freedom varies from one configuration to another: for example, the AIC, the Bayesian Information Criterion (BIC; Schwarz, 1978), the Geometric Information Criterion (GIC; Kanatani, 1998), or the Stochastic Information Complexity (SIC; Rissanen, 1989). However, in our pulsar observations declinations are not uniformly sampled and only different declination values are tested, where is the number of pulsars observed, whereas hour angles are sampled uniformly and the number of hour angles tested is much larger than . As a consequence, strong asymmetries exist between the two dimensions, in terms of the information available to constrain model parameters. The AIC, BIC and GIC metrics, that are in principle designed to penalize model complexity by accounting for the varying numbers of degrees of freedom in the compared models, cannot account for such strong asymmetries in the sampled dimensions. The SIC metric, which we also tested, uses the logarithm of the determinant of the Hessian metrix to penalize model complexity. However, we found that the Hessian tends to be ill-conditioned in the cases where the polynomial order in declination approaches the number of pulsars available to constrain any model of variation along this dimension. Finally, additional complexity emerged from the fact that pcm analyses tended to successfully fit the available data in fewer frequency channels when polynomial orders increased. With all the above-mentioned caveats in mind, we chose to adopt the simple AIC metric in our tests, but dividing AIC scores by the number of frequency channels with successful fits, to form the “quasi-AIC” metric. In future work we will further investigate the issue of the choice of metric for comparing the results from these analyses. We also note that in addition to qAIC scores, we used pulsar observation S/N values as another metric for comparing the different configurations, as is presented later in this section.
@@
In Fig. 4, we show the best-fit calibration parameters and Stokes parameters of the noise diode from the 0/0 analysis of the rotating horn observations made between MJDs 59368 and 59373, and those from the analysis of the observation of J07422822 on MJD 56368. In general the results from the two analyses are very similar. In some cases, such as those of the differential gain or the ellipticity parameters and (assumed to be equal), some small differences are found. Nevertheless, best-fit parameters from the 0/0 analysis tend to present smoother variations across the bandwidth than those from the analysis of J07422822 and preceding noise diode observation, with less pronounced variability and fewer outliers (see the best-fit values for the orientation parameter for a particularly clear example). This suggests that the instrumental response inferred from the 0/0 analysis better represents the actual response than that inferred from the analysis of J07422822, and highlights the usefulness of jointly analyzing multiple pulsar observations at once.
@@
Most other analyses resulted in higher qAIC scores than the simplest 0/0 model, with the exceptions of the 0/1, 1/0 and 1/1 configurations which gave very similar qAIC scores. In order to discriminate between these models, we conducted two other series of rotating horn observations of pulsars at different declinations, between MJDs 60240 (23 October 2023) and 60242 (25 October 2023), and on MJD 60461 (31 May 2024). These second and third observation sessions are described in Table 1. The pulsars observed during these two new sessions were the same as in the first, with the exception of PSR J0358+5413 (B0355+54, Manchester et al., 1972) which was added to the target list for the second session. Beyond the fact of simply taking more data, we took the opportunity of these new observing sessions to observe over longer durations than what is typically done with the NRT, and thereby observe over wider HA ranges, close to the limit that can explored with this telescope. With these longer observations, we wanted to increase our sensitivity to potential dependencies of calibration parameters on HA.
@@
We carried out pcm analyses for each of these three observation sessions, testing the same combinations of 2D polynomial orders as described above. The results from these combined analyses are summarized in Table 2. In this table, the “Total AIC” column gives the sum of the AIC scores from the three observation sessions, the “Total nchan” columns gives the total number of frequency channels in which the fit was successful (since the NUPPI backend splits the 512 MHz of bandwidth into 128 channels of 4 MHz each, the maximum possible value is ), and the “qAIC” column gives the quasi-AIC scores, i.e., the total AIC scores divided by the total number of frequency channels with a succesful fit. As can be seen from the table, for configurations with polynomial orders in declination larger than three, the total numbers of frequency channels with a successful fit are smaller than under the simplest configurations, and even much smaller in some cases (see, e.g., the results for the 6/3 analysis), indicating unstable fits. Quasi-AIC scores are also higher, with the exception of the 4/3 configuration which gave the lowest qAIC score among all analyses. In this analysis, however, successful fits were found in much fewer frequency channels than for the simplest analyses, so that we did not consider the 4/3 configuration further. The simplest configurations with polynomial orders in declination of up to two are found to produce the best results, both in terms of qAIC score and number of frequency channels with successful fits.
@@
Taken at face value, Table 2 suggests that 0/1, 0/2, 0/3 and 1/1 configurations produce better results than the simplest model, the 0/0 configuration. To test whether these seemingly better configurations also constitute better calibration solutions for normal mode observations (i.e., regular NRT observations of pulsars with no horn rotation), we analyzed selections of NUPPI observations of MSPs, calibrating the observations with the parameters found by pcm for each of the three sessions and configurations. The selections consisted in MSP observations conducted within 30 days of MJDs 59370, 60240, and 60461, i.e., close in time to the three special mode observation sessions presented in Table 1. We selected normal mode MSP observations with S/N values higher than 100, to limit the sample of observations to analyze444The S/N values considered here were those obtained after calibrating the NUPPI data with the method presented in Guillemot et al. (2023), i.e., using calibration solutions derived from rotating horn observations of PSR J07422822.. In this work, S/N values were calculated using Eq. (7.1) of Lorimer & Kramer (2004):
@@
Here denotes the amplitude of the -th profile bin, and represent the off-pulse standard deviation and mean amplitude, and is the equivalent width, expressed in number of profile bins, of a top-hat pulse with the same area and peak height as the profile. After the above-mentioned selection cuts, we obtained datasets comprising 52, 55 and 57 MSP observations, respectively.
@@
In Fig. 5, we compare the S/N values of the total intensity pulse profiles for the three MSP datasets under different calibration methods. In the left panel we compare S/N values as obtained when calibrating MSP observations with the 0/0 model, and those obtained using the IFA calibration model. The improved calibration produces increased S/N values in more than 75% of the considered observations, in line with the results of Guillemot et al. (2023). The middle and right panels show comparisons of the 0/0 model S/N values with those obtained with the 0/1 and 1/1 calibration models. S/N values obtained with the 0/0 and 0/1 models are very similar, with differences very close to 0 in most cases. Yet, more than 60% of the S/N values are higher with the 0/0 model than with the 0/1 model. As can be seen from the right panel, S/N values obtained when calibrating the data with the 1/1 model (i.e., with differential gain and phase values that depend on declination) are degraded in almost 80% of cases, with, in particular, large differences at high declination values. We also calibrated the same datasets using 0/2 and 0/3 models, finding that these models degraded S/N values in 79% and 78% of the cases, respectively. We therefore find that the 0/0 model, in which the differential gain and phase parameters do not vary with declination and hour angle, produce better results than the other models on normal mode MSP observations, in terms of S/N values. This suggests that the better qAIC values obtained for the 0/1, 0/2, 0/3 and 1/1 configurations (see Table 2) resulted from overfitting of the data, and not from actual improvements of the models. Finally, we repeated the same analyses as those presented earlier in this section, this time allowing the feed orientation parameter and the feed ellipticities to vary with hour angle and declination using the same parameterization as presented in Eq. 3, and fitting constant values for the differential gain and phase parameters. The conclusions from this analysis were similar: better qAIC scores were found for a few of the tested configurations (the same configurations as in Table 2) than for the 0/0 one. However, the application of these calibration solutions to normal-mode observations of MSPs made within 30 days of MJDs 59370, 60240, and 60461 again led to degraded S/N values. We therefore take our results to mean that the polarimetric response of the NRT does not seem to depend on hour angle and declination.
@@ 3 Improving polarization calibration over the entire NUPPI dataset
The first analysis step was thus to construct a high S/N well-calibrated polarimetric profile, of a particular pulsar. Ideally, the chosen pulsar should be bright enough for METM to be able to determine accurate calibration solutions for the individual 4 MHz frequency channels. Tests we conducted showed that this condition requires individual observations to have S/N values of several thousands at least. In addition, there need to be numerous observations of the desired pulsar, between August 2011 and the end of the considered NUPPI dataset. Two pulsars stand out, in the case of the NUPPI data archive: J0953+0755 (B0950+08), and J1136+1551 (B1133+16; Pilkington et al., 1968). Among these two objects, J0953+0755 has the highest flux density at 1.4 GHz: against mJy (Jankowski et al., 2018), and is therefore typically detected with higher S/N values with the NRT. J1136+1551, however, has been observed much more often than J0953+0755 with NUPPI: almost four times more often, so that the density of J1136+1551 observations is larger. Given that both pulsars appeared to be suitable choices for the METM analysis, we conducted analysis tests using both of them.
@@
We selected individual 1.4 GHz NUPPI observations of J0953+0755 and J1136+1551, conducted close in time to the first session described in Table 1 and with high S/N values. The selection led us to choose the MJD 59372 observation of J0953+0755 listed in Table 1, and a normal mode observation of J1136+1551 conducted on MJD 59376 (11 June 2021). The two observations were calibrated using a modified version of the best-fit calibration solution obtained from the 0/0 analysis of the rotating horn observations made between 59368 and 59373. This modified version consisted in a version of the best-fit calibration solution in which parameters for the frequency channels with no data (e.g., due to RFI mitigation) were linearly interpolated using data from the neighboring channels. The selected and calibrated NUPPI observations were integrated in time, and the full frequency resolution of 128 channels was kept. Both observations had high S/N values: 14000 for J0953+0755 and 4480 for J1136+1551.
@@
All other 1.4 GHz NUPPI observations of J0953+0755 and J1136+1551 (and preceding observations of the noise diode) were then compared to the above-described reference polarimetric profiles, using the METM method of pcm. In addition to determining the gain , differential gain , differential phase and the Stokes parameters of the reference noise diode, pcm fits for four additional parameters: the difference in the receiver ellipticities, , the sum of the ellipticities, , the difference in the receiver orientations, and the sum of the orientations, . Inspection of the analysis results for the individual observations of J0953+0755 and J1136+1551 enabled us to observe that the best-fit calibration parameters as determined from analyses of J1136+1551 observations displayed very strong variability across the bandwidth, strong variability from epoch to epoch, and large numbers of outlier values. On the other hand, best-fit results from the analysis of J0953+0755 observations displayed much smoother variations, and with small numbers of outlier values. The much better fits obtained in the case of J0953+0755 led us to choose this pulsar for the subsequent analyses, and discard J1136+1551.
@@
In order to calibrate a normal mode NUPPI observation of a pulsar at a given epoch, one possibility is to use the calibration parameters derived using the METM method from the J0953+0755 observation made closest to that epoch. However, in further investigating the results for individual J0953+0755 observations, we noticed that the frequency variations in the best-fit parameters were consistent over time segments. More specifically, in each of these time segments, the individual parameters appeared to be scaled and/or offset versions of common functions, to be determined. Instead of calibrating pulsar observations using the calibration parameters derived from the J0953+0755 observation conducted closest in time, we thus chose to explore a different strategy: determine the boundaries of the above-mentioned time segments, determine the “archetypal” functions describing the frequency variations of the different calibration parameters over those time segments, and eventually use this information to find the calibration parameters at the considered epoch.
@@
As mentioned above, best-fit results from the analysis of J0953+0755 were more suitable for the analysis presented here than those obtained from the analysis of J1136+1551 observations. Nevertheless, the larger density of J1136+1551 observations made the data for the latter pulsar more appropriate for finding the boundaries of the different time segments. In Fig. 6 we show three of the calibration parameters determined with the METM method as a function of time, for PSR J1136+1551, and for the 64th 4 MHz frequency channel. As can be noted from this figure, NUPPI observations of J1136+1551 have been conducted regularly over the entire dataset, with a small number of gaps that generally coincide with time intervals over which the NRT was inactive (see Sect. 3.1 of Guillemot et al., 2023, for a list of the main interruptions between 2011 and 2023). Vertical black lines indicate J0953+0755 observations, also conducted regularly since August 2011, but with lower cadence. We visually inspected METM results for J1136+1551, searching for time discontinuities in the frequency variations of best-fit parameters. In Fig. 7 we illustrate, with the example of the absolute gain parameter , the variations seen from one time interval to another in the evolution of calibration parameters with frequency. This step enabled us to determine the boundaries of the above-mentioned time segments, indicated as dashed vertical lines in the figure.
@@
In the following step, for each time segment we selected a single observation and used the calibration parameters for this observation as initial guesses of the “archetype” functions for the corresponding time segment. The single observations we chose as starting points had good fits and small numbers of excised frequency channels. The frequency evolution of the parameters for the chosen observations were spline smoothed to form the initial guesses of the archetypes. Then, the parameters for the other observations in the time segments were fitted to their corresponding archetypes. The gain parameter was fitted using the following function:
@@
Here represents the value of the archetype function at frequency , is the value of the parameter to be fitted and is a scale parameter. The and parameters, found to be close to zero in most channels and observations, were fitted to the function
@@
where is an offset parameter. Finally, the other parameters (, , , , and Stokes parameters , , and for the reference noise diode) were fitted as
@@
In the following analysis step we refined the archetypes for each parameter and time segment, by forming weighted averages of the calibration parameters within each segment, correcting for the scale and offset parameters and for each observation determined from the previous step. This analysis step enabled us to obtain improved parameter archetypes, that in some cases included values for more frequency channels, i.e., the frequency channels that were absent in the chosen starting points but present in other observations from the same time segment. The improved parameter archetypes were then used to determine new estimates of the scale and offset parameters and using Eq. 7 to 9, for each calibration parameter and time segment. In Fig. 8 we plot the values of the calibration parameters as a function of frequency for two of the time segments. Parameter values were multiplied and/or offset using the and factors determined from this iteration of the analysis. The scaled and offset calibration parameters match the corresponding archetypes well, validating the construction of the archetypes and the measurement of the and parameters for each observation. Fig. 8 illustrates the fact that calibration parameters have common frequency evolution properties in each time segment, and that the frequency dependence of the parameters change significantly in other time segments.